# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

### v1

@@ Abstract
Model editing techniques are essential for efficiently updating knowledge in large language models (LLMs). However, the effectiveness of existing approaches degrades in massive editing scenarios, particularly when evaluated with practical metrics or in context-rich settings. We attribute these failures to embedding collisions among knowledge items, which undermine editing reliability at scale. To address this, we propose NAMET (Noise-aware Model Editing in Transformers), a simple yet effective method that introduces noise during memory extraction via a one-line modification to MEMIT. Extensive experiments across six LLMs and three datasets demonstrate that NAMET consistently outperforms existing methods when editing thousands of facts.
@@ 1 Introduction
Large language models (LLMs) are increasingly employed as search engines and chatbots, as they excel at retrieving knowledge to answer user queries [4, 31, 34, 3]. However, they are prone to spreading misinformation about frequently updated topics due to outdated training data [32, 11, 33]. To address this issue, retraining or fine-tuning models for partial knowledge updates is proposed [1, 28], albeit with prohibitively expensive overhead. In contrast, recent advances in model editing (ME) techniques [20, 6] enable massive editing of thousands of factual associations concurrently at minimal data and computational cost, thereby rendering real-time knowledge updates feasible.
Figure 1: (LEFT) Comparison of editing efficacy under both metrics on the CounterFact dataset, with and without prefixes. The editing performance is evaluated using MEMIT [20], PMET [15], and NAMET. The loose metric compares the next-token probability of the target and original objects, while the practical metric checks whether the target object is explicitly generated. (RIGHT) Distribution of edited knowledge representations extracted using MEMIT and NAMET.
@@
Despite the success of existing massive ME techniques, we observe that their reported performance is often overestimated due to overly loose evaluation metrics. In particular, most prior works assess editing quality by checking whether the model is more likely to generate the following tokens as the target object than the original one, whereas neglecting to evaluate whether the model’s output is consistent with the target object [20, 6]. Therefore, we advocate a “practical metric”, which measures the proportion of cases in which the edited model retrieves the target object and explicitly generates related output. This metric provides a more accurate reflection of real-world usage, as will be shown in our evaluation setting (see Section 6). Considering the left of Figure 1, the loose metric does not require the edited model to explicitly generate the target object when editing "Jeep Commander is created by Honda." Under the loose metric, an edited model that outputs "Jeep Commander is created by the American company Jeep" may still be considered a successful edit, whereas the strict metric would classify it as a failure. Under such evaluation criteria, these methods fail to maintain performance in massive editing settings, as shown on the left of Figure 1. Moreover, existing approaches exhibit substantial degradation when edited knowledge is preceded by prefixes [17], which is common in realistic question-answering tasks [24, 26]. This lack of robustness to contextual variations undermines their practicality in real-world applications.
To achieve robust massive editing under practical settings, we propose NAMET (Noise-Aware Model Editing in Transformers), which outperforms existing approaches under stricter evaluation criteria and exhibits strong robustness to contextual variations. We first identify “embedding collisions”, which reflect the distinctiveness of hidden representations among edited facts, as a primary cause of performance degradation in both massive editing and context-rich scenarios. NAMET mitigates this issue by introducing noise during target memory optimization via a simple one-line modification to MEMIT. As shown on the right of Figure 1, the introduced noise reduces the occurrence of embedding collisions, thereby enabling a greater number of edits to be performed simultaneously.
@@
In this paper, extensive experiments on six LLMs demonstrate that NAMET consistently outperforms existing methods under more rigorous scenarios across the CounterFact, ZsRE, and Wiki-recent datasets. Specifically, on the CounterFact dataset, NAMET achieves up to a 15% improvement in editing efficacy over existing methods, and maintains over 90% editing efficacy across all evaluated LLMs when editing 10,000 facts. Furthermore, NAMET consistently maintains high editing accuracy even when edited items are preceded by up to 200-token prefixes, highlighting its robustness in realistic, context-rich settings, including chatbots and long-context QA tasks.
@@ 2 Related Work
Model Editing. Existing ME techniques can be classified into auxiliary-based [10, 22, 38, 36, 21] and location-based methods [19, 20, 15]. Auxiliary-based ME techniques preserve the original parameters, and introduce additional information to edit knowledge. SERAC [22] requires extra memory to store new edits and learn to reason over them to manipulate the model’s output. Location-based methods directly modify model parameters to edit knowledge without requiring any additional information. These methods assume that factual associations are stored in the feed-forward networks (FFNs) of the LLMs [8, 7, 5]. Building on these, ROME [19] first gains insights on the specific location of the knowledge through causal analysis. It proceeds to directly modify critical MLP layers to update factual associations. MEMIT [20] builds upon ROME to enable massive editing of thousands of facts concurrently. AlphaEdit [6] focuses on sequential editing, aiming to preserve both previously edited knowledge and the general capabilities of the LLM during successive edits.
Massive Editing. In practical applications, ME techniques may aim to update a model with hundreds or even thousands of facts simultaneously in order to keep up with the constantly evolving knowledge [13, 9]. However, auxiliary-based methods are limited in scalability, typically supporting only a few edits at a time [22]. In contrast, location-based methods are more scalable for massive editing. MEMIT [20] scales to edit 10,000 facts concurrently, and PMET [15] further improves performance by incorporating attention layers when updating the parameters of the FFNs. Despite their effectiveness and scalability, these methods have been shown to be fragile when handling prefixes during evaluation, which is a common scenario in real-world applications [17, 35, 18]. Moreover, we observe that their performance in massive editing is overestimated due to the loose metric. In this work, we propose NAMET, which achieves superior performance in massive editing under practical evaluation metrics, while also exhibiting greater robustness against long prefixes.
@@ 3 Preliminary: Editing Memory in LLMs
Previous works have shown that a pre-trained LLM has memorized many factual associations [23, 12, 25, 27]. These stored facts could be edited by modifying the MLP layers within FFN modules, based on the assumption that knowledge is stored in them in the form of key-value pairs [8, 7].
@@
In Figure 2, the MLP layer $W_{out}^{l}$ within FFN associates keys $k_{t}^{l}(x)=\sigma(W^{l}_{in}\gamma(h^{l-1}_{t}(x)))$ with memories $m^{l}_{t}(x)$ for the fact $x$. Given the critical mediating role of MLP layers in storing facts, Meng et al. [19] shows that it is sufficient to update $W_{out}^{l}$ to edit stored facts. We then optimize $W_{out}^{l}$ (abbreviated as $W_{1}$) as follows:
$W_{1}\triangleq\underset{\hat{W}}{\arg\min}(\sum_{i=1}^{N_{t}}||\hat{W}k_{i}^{t}-m_{i}^{t}||^{2}+\sum_{j=1}^{N_{p}}||\hat{W}k_{j}^{p}-m_{j}^{p}||^{2})$ (1)
Here, $k_{i}^{t}$ and $k_{j}^{p}$ denote the encoded subject representations for individual target and preserved fact $i$ and $j$, respectively, while $m_{i}^{t}$ and $m_{j}^{p}$ represent their corresponding memory vectors. We stack the keys and memories of totally $N_{t}$ target knowledge into matrices as $K_{t}=[k_{1}^{t}\ |\ k_{2}^{t}\ |\ \dots\ |\ k_{N_{t}}^{t}]$ and $M_{t}=[m_{1}^{t}\ |\ m_{2}^{t}\ |\ \dots\ |\ m_{N_{p}}^{t}]$. Similarly, we construct $K_{p}$ and $M_{p}$ for $N_{p}$ preserved facts. The objective in Equation 1 can then be optimized by solving the normal equations [20]:
@@
where we expand $W_{1}$ into $W_{0}+\Delta$. $W_{0}$ denotes the original (unedited) parameters that associate preserved keys with their memory representations. The final update to $W_{out}^{l}$ can be computed by multiplying both sides of Equation 2 by $\begin{bmatrix}K_{p}&K_{t}\end{bmatrix}^{T}$, and subtracting Equation 3 from Equation 2[20]:
@@
where $R=M_{t}-W_{0}K_{t}$ denotes new relations’ residual with respect to the original weights, which can also be written as $[r_{1}^{t}\ |\ r_{2}^{t}\ |\ \ldots\ |\ r_{N_{t}}^{t}]$. Since the pretraining data of the original model is not accessible, we approximate $C_{p}$ using a set of randomly sampled inputs from public datasets:
@@
Despite their effectiveness, existing methods struggle when concurrently editing a large number of facts or when the edited knowledge is preceded by prefixes. Here, we termed $r_{i}k_{i}^{T}$ the residual-key pair, and $k_{i}k_{i}^{T}$ the key-key pair for target fact $i$ [^1] We omit superscript $t$ for clarity. We then analyze these challenges through the lens of embedding collision of residual-key (abbreviated as r-k) and key-key (abbreviated as k-k) pairs.
@@ 4 Motivation
In this section, we investigate the underlying causes of the challenges associated with massive editing, as illustrated in Figure 1. Specifically, we examine the decline in editing performance as the number of edited facts increases, and the specific failure cases where edits are preceded by prefixes. Our theoretical and empirical analyses suggest these issues stem from embedding collisions, reflecting the insufficient distinctiveness between r-k and k-k pairs in the embedding space.
4.1 Theoretical Analysis of Embedding Collision
We observe that by expanding $K_{t}$ and $R$ in Equation 4, the update equation can be reformulated as:
@@ 4.1 Theoretical Analysis of Embedding Collision
where the update $\Delta$ is determined by the aggregated r-k and k-k embeddings across all edited facts. As the number of edits increases, solving Equation 6 is more likely to cause embedding collisions between the r-k and k-k embeddings associated with a specific fact with other facts. This eventually leads to information interference and degraded editing performance.
@@
To formalize the concept of embedding collision, we define two key requirements for the desired update $\Delta$: (1) The update should preserve the existing knowledge, expressed as $\Delta C_{p}=0$. (2) The update should ensure lossless reconstruction for each individual fact, formulated as $\Delta k_{i}k_{i}^{T}=r_{i}k_{i}^{T}$, where $\Delta$ is computed while considering all target facts. For notational simplicity, we define $S_{i}=k_{i}k_{i}^{T}$ and $T_{i}=r_{i}k_{i}^{T}$. Incorporating (1), an ideal $\Delta$ that meets (2) implies:
$\Delta\sum_{i=1}^{N_{t}}S_{i}=\sum_{i=1}^{N_{t}}T_{i}\quad\rightarrow\quad\Delta S_{i}=T_{i}\quad\text{for }i=1,2,\ldots,N_{t}$ (7)
@@
However, the validity of Equation 7 is intuitively affected by the degree of orthogonality among the embeddings of different facts. We then define this orthogonality as embedding collision:
Definition 1 (Embedding Collision). Given two knowledge items $i$ and $j$, and an embedding matrix $E\in\{S,T\}$, the embedding collision between $i$ and $j$ is the orthogonality between their respective embedding vectors $E_{i}$ and $E_{j}$, which could be measured by their cosine similarity $cos(E_{i},E_{j})$.
@@
We now formalize the connection between embedding collision and the editing performance of a specific knowledge item $i$ under massive editing. Specifically, we quantify the degree to which Equation 7 is established by analyzing the reconstruction loss $RL_{i}=\Delta S_{i}-T_{i}$ for each knowledge item. This relationship is formalized in the following theorem:
Theorem 1. The reconstruction loss for the $i$-th knowledge item is given by:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i}$ (8)
where the first term, $T_{i}S^{+}S_{i}-T_{i}$, is amplified by the cosine similarity (collision) between $S_{i}$ and other $S_{j}$ embeddings for $j\neq i$. The second term, $\sum_{j\neq i}T_{j}S^{+}S_{i}$, is jointly amplified by the collision between $T_{i}$ and $T_{j}$, as well as between $S_{i}$ and $S_{j}$ for $j\neq i$.
@@
This result demonstrates how embedding collisions impact the editing performance of individual knowledge items under massive editing. Specifically, stronger collision among knowledge items leads to increased individual reconstruction loss, ultimately reducing the overall effectiveness of massive editing. The complete proof is provided in Appendix A.
4.2 Empirical Study of Embedding Collision
Motivated by the above analysis, we thus hypothesize that fostering more “isolated” r-k and k-k representations across knowledge items would enhance capacity for massive concurrent editing. Such isolation is also vital for robustness against long prefixes, where collisions typically exacerbate performance degradation. This intuition motivates two hypotheses below, which we propose and empirically verify, to explain failures in existing ME techniques, followed by our introduction of a simple yet effective method to address these challenges.
Hypothesis 1. As the number of edited knowledge items increases, the r-k and k-k pairs associated with a specific item are more likely to collide with those of other items in the embedding space, leading to degraded editing performance for that specific item.
Hypothesis 2. Long prefixes exacerbate embedding collision issues under massive editing, leading to degraded editing performance when edited facts are evaluated with descriptive prefixes.
We proceed to verify the above hypotheses through carefully designed experiments.
[…]
LLaMA2-7B 200 98.5 0.09 66.66% 0.19 66.66% 21.79 100.00% 100.00%
500 90.0 0.20 74.00% 0.41 78.00% 21.21 74.00% 92.00%
1000 86.8 0.40 69.70% 0.72 72.73% 20.16 81.82% 95.45%
Deepseek-7B 200 99.5 0.11 100.00% 0.21 100.00% 4.35 100.00% 100.00%
500 98.6 0.27 85.71% 0.60 100.00% 4.96 100.00% 100.00%
1000 97.8 0.52 81.81% 1.03 100.00% 7.03 86.36% 100.00%
@@ 4.2 Empirical Study of Embedding Collision
Empirical Verification of Hypothesis 1. We separately modify 200, 500, and 1,000 knowledge items in LLaMA2-7B [30] and Deepseek-7B [3] using MEMIT [20]. These facts are sampled from the CounterFact dataset [19]. We evaluate the editing accuracy of these items where no prefix precedes the edited query. We measure the embedding collision using the collision score, which is defined as follows:
Definition 2 (Collision Score). Given a knowledge item $i$, and an embedding matrix $E\in\{S,T,\lambda\}$, its collision score $V^{E}_{i}$ is computed as the average cosine similarity between its embedding and those of all edited facts:
$V^{E}_{i}=\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\cos(E_{i},E_{j})$ (9)
For each knowledge item, we compute the collision score for their r-k ($T$) and k-k ($S$) embeddings. To incorporate the influence of $C_{p}$, we also calculate the collision score of the k-k-$\lambda$ pairs, defined as $\frac{1}{n}C_{p}+k_{i}k_{i}^{T}$ for each item. We aggregate the collision scores across all items to quantify the overall degree of collision. To further validate Hypothesis 1, we identify the items whose target objects fail to be retrieved during evaluation. We then examine whether their corresponding r-k, k-k, and k-k-$\lambda$ collision scores rank in the top 50% among all edited items.
As shown in Table 1, the decline in overall editing accuracy for both models is correlated with an increase in embedding collision when more facts are edited. Specifically, for LLaMA2-7B, editing accuracy decreases from 98.5% to 86.8% as the number of edited facts rises from 200 to 1,000. Correspondingly, the collision scores increase from 0.09 to 0.40 for r-k pairs and from 0.19 to 0.72 for k-k pairs. At the individual level, editing performance is similarly impacted by collision degree. Approximately 70% of r-k and k-k pairs associated with failed edits rank within the top 50% of collision scores. Incorporating $C_{p}$ further amplifies this effect, raising the proportion of high-collision items among failure cases to 100% for 200 edited items and 81.82% for 1,000 items. Overall, more than 90% of all failed edits exhibit at least one of the three collision scores (r-k, k-k, or k-k-$\lambda$) ranking within the top 50% of the collision distribution.
We also observe that, despite exhibiting higher collision scores for both r-k and k-k pairs compared to LLaMA2-7B, Deepseek-7B still achieves superior editing accuracy. This discrepancy does not contradict our hypothesis, as comparisons between embeddings across models may be influenced by variations in inherent model settings. For a more precise comparison, we extract the principal components of r-k and k-k pairs from the same 100 items in both LLaMA2-7B and Deepseek-7B, and visualize them in Figure 3. The r-k and k-k embeddings in Deepseek-7B are significantly more dispersed than those in LLaMA2-7B, suggesting that Deepseek-7B’s embeddings are less prone to collisions. This observation aligns with its superior editing performance, further supporting our hypothesis.
Figure 3: PCA visualization of r-k and k-k embeddings for 100 edited facts at layer 3 in LLaMA2-7B and Deepseek-7B, using MEMIT.

### v4

# EAMET: Robust Massive Model Editing via Embedding Alignment Optimization

@@ Abstract
Model editing techniques are essential for efficiently updating knowledge in large language models (LLMs). However, the effectiveness of existing approaches degrades in massive editing scenarios, particularly when evaluated with practical metrics. Their robustness is also limited in context-rich settings or when editing multiple facts of the same subject simultaneously. We attribute these failures to the embedding misalignment among knowledge items, which undermines editing reliability at scale. To address this, we propose EAMET (Embedding Alignment Model Editing in Transformers), which addresses this issue by aligning the space of key and residual embeddings. Extensive experiments across six LLMs and three datasets demonstrate that EAMET consistently outperforms existing methods, achieving about 90% editing efficacy when editing 10k facts. Codes and datasets are publicly available at https://github.com/ybdai7/EAMET-massive-editing.
@@ 1 Introduction
Large language models (LLMs) are increasingly employed as search engines and chatbots, as they excel at retrieving knowledge to answer user queries (Brown et al., 2020; Touvron et al., 2023b; Yang et al., 2024a; Bi et al., 2024). However, they are prone to spreading misinformation about frequently updated topics due to outdated training data (Vykopal et al., 2023; Huang et al., 2025; Xu et al., 2024). To address this issue, retraining or fine-tuning models for partial knowledge updates is proposed (Achiam et al., 2023; Team et al., 2025), albeit with prohibitively expensive overhead. In contrast, recent advances in locate-then-edit model editing (ME) techniques (Meng et al., 2023; Fang et al., 2024) enable massive editing of thousands of factual associations concurrently at minimal data and computational cost, thereby rendering real-time knowledge updates feasible.
Figure 1: Illustration of current methods and our proposed EAMET in evaluating massive editing. Here, “[PREFIX] Sentence” and “{Sentence $\mid$ $s_{i}=\text{Jeep Commander}$}” denote the scenarios where the edited knowledge is preceded by prefixes and where multiple facts share the same subject, respectively.
@@
Despite the success of existing massive ME techniques, we observe that their effectiveness is often overestimated due to overly loose evaluation metrics. In particular, most prior works assess editing quality by checking whether the model is more likely to generate the following tokens as the target object than the original one, whereas neglecting to evaluate whether the model’s output is consistent with the target object (Meng et al., 2023; Fang et al., 2024). Therefore, we advocate a “practical metric”, which measures the proportion of cases in which the edited model retrieves the target object and explicitly generates related output. This metric provides a more accurate reflection of real-world usage, as will be shown in our evaluation setting (see Section˜6). Under such evaluation criteria, existing methods fail to maintain their performance.
Moreover, existing methods exhibit limited robustness in realistic settings. We highlight two representative scenarios: (i) their performance substantially degrades when edited knowledge is preceded by prefixes (Li et al., 2024), a common phenomenon in practical question-answering tasks (Pramanick et al., 2024; Romero et al., 2024); and (ii) they fail to preserve accuracy when editing multiple facts associated with the same subject, where performance drops markedly. Such lack of robustness in massive editing scenarios undermines their applicability to real-world use cases.
To analyze the limitations of existing methods, we first identify “embedding misalignment”, which reflects the structural inconsistency between key and residual embedding spaces, as a primary factor underlying the decline in both effectiveness and robustness during massive editing. Such misalignment leads to information loss for individual knowledge updates. In particular, when parameters are updated jointly from a batch of edited knowledge items, they fail to accurately reconstruct an individual factual association. This information loss becomes more severe as the number of edited items increases.
To achieve effective and robust massive editing under practical settings, we thus propose EAMET (Embedding Alignment Model Editing in Transformers), which outperforms existing approaches under stricter evaluation criteria and exhibits strong robustness in two described scenarios. EAMET addresses embedding misalignment by progressively preserving optimized residual embeddings and aligning them with the key embedding space, ensuring consistency throughout the editing process.
@@
In this paper, we conduct extensive experiments on six LLMs, showing that EAMET consistently surpasses existing methods under rigorous settings across the CounterFact, ZsRE, and Wiki-recent datasets. EAMET maintains about 90% editing efficacy across all evaluated models and outperforms baselines by an average of 14% and 8%, with gains of up to 37% and 15% on CounterFact and ZsRE when editing 10k facts. Moreover, EAMET sustains high accuracy even when edited items are preceded by prefixes of up to 200 tokens or involve multiple facts associated with the same subject. This demonstrates EAMET’s robustness in realistic and context-rich settings, including chatbots and long-context QA tasks.
@@ 2 Related Work
Model Editing. Existing ME techniques can be classified into auxiliary-based (Hartvigsen et al., 2023; Mitchell et al., 2022b; Zheng et al., 2023; Yu et al., 2024; Mitchell et al., 2022a) and location-based methods (Meng et al., 2022; 2023; Li et al., 2025). Auxiliary-based ME techniques preserve the original parameters, and introduce additional information to edit knowledge. SERAC (Mitchell et al., 2022b) requires extra memory to store new edits and learn to reason over them to manipulate the model’s output. Location-based methods directly modify model parameters to edit knowledge without requiring any additional information. These methods assume that factual associations are stored in the feed-forward networks (FFNs) of the LLMs (Geva et al., 2021; 2022; Dai et al., 2022). Building on these, ROME (Meng et al., 2022) first gains insights on the specific location of the knowledge through causal analysis. It proceeds to directly modify critical MLP layers to update factual associations. MEMIT (Meng et al., 2023) builds upon ROME to enable massive editing of thousands of facts concurrently. AlphaEdit (Fang et al., 2024) focuses on sequential editing, aiming to preserve both previously edited knowledge and the general capabilities of the LLM during successive edits.
Massive Editing. In practical applications, ME techniques may aim to update a model with hundreds or even thousands of facts simultaneously in order to keep up with the constantly evolving knowledge (Ju et al., 2024; Gu et al., 2024). However, auxiliary-based methods are usually limited in scalability, typically supporting only a few edits at a time (Mitchell et al., 2022b). In contrast, location-based methods are more scalable for massive editing. MEMIT (Meng et al., 2023) scales to edit 10,000 facts concurrently, and PMET (Li et al., 2025) further improves performance by incorporating attention layers when updating the parameters of the FFNs. Despite their effectiveness and scalability, these methods have been shown to be fragile when handling prefixes or multiple facts with the same subject during evaluation, which is a common scenario in real-world applications (Li et al., 2024; Yang et al., 2024b; Ma et al., 2024). Moreover, we observe that their performance in massive editing is overestimated due to the loose metric. In this work, we propose EAMET, which achieves superior performance in massive editing under practical evaluation metrics, while also exhibiting greater robustness against long prefixes and multiple facts with the same subject.
@@ 3 Preliminary: Editing Memory in LLMs
Previous works have shown that a pre-trained LLM has memorized many factual associations (Petroni et al., 2019; Jiang et al., 2020; Roberts et al., 2020; Shin et al., 2020). These stored facts could be edited by modifying the MLP layers within FFN modules, based on the assumption that knowledge is stored in them in the form of key-value pairs (Geva et al., 2021; 2022).
@@
In Figure˜2, the MLP layer $W_{out}^{l}$ within FFN associates keys $k_{t}^{l}(x)=\sigma(W^{l}_{in}\gamma(h^{l-1}_{t}(x)))$ with memories $m^{l}_{t}(x)$ for the fact $x$. Given the critical mediating role of MLP layers in storing facts, Meng et al. (Meng et al., 2022) shows that it is sufficient to update $W_{out}^{l}$ to edit stored facts. We then optimize $W_{out}^{l}$ (abbreviated as $W_{1}$) as follows:
[…]
Here, $k_{i}^{t}$ and $k_{j}^{p}$ denote the encoded subject representations for individual target and preserved fact $i$ and $j$, respectively, while $m_{i}^{t}$ and $m_{j}^{p}$ represent their corresponding memory vectors. We stack the keys and memories of totally $N_{t}$ target knowledge into matrices as $K_{t}=[k_{1}^{t}\ |\ k_{2}^{t}\ |\ \dots\ |\ k_{N_{t}}^{t}]$ and $M_{t}=[m_{1}^{t}\ |\ m_{2}^{t}\ |\ \dots\ |\ m_{N_{p}}^{t}]$. Similarly, we construct $K_{p}$ and $M_{p}$ for $N_{p}$ preserved facts. The objective in Equation˜1 can then be optimized by solving the normal equations (Meng et al., 2023):
@@
where we expand $W_{1}$ into $W_{0}+\Delta$. $W_{0}$ denotes the original (unedited) parameters that associate preserved keys with their memory representations. The final update to $W_{out}^{l}$ can be computed by multiplying both sides of Equation˜2 by $\begin{bmatrix}K_{p}&K_{t}\end{bmatrix}^{T}$, and subtracting Equation˜3 from Equation˜2(Meng et al., 2023):
@@
where $R=M_{t}-W_{0}K_{t}$ denotes new relations’ residual with respect to the original weights, which can also be written as $r_{1}^{t}\ |\ r_{2}^{t}\ |\ \ldots\ |\ r_{N_{t}}^{t}]$. Since the pretraining data of the original model is not accessible, we approximate $C_{p}$ using a set of randomly sampled inputs from public datasets:
@@ 4 Motivation
In this section, we investigate the root causes of the challenges associated with effective and robust massive editing, as illustrated in Figure˜1. In particular, we analyze the decline in editing performance as the number of edited facts increases. Our theoretical and empirical results indicate that these issues arise from misalignment between key and residual embeddings. We further examine robustness in two representative scenarios: (i) edits preceded by long prefixes, and (ii) edits applied to multiple facts sharing the same subject.
4.1 Embedding Misalignment in Effective Massive Editing
Theoretical Analysis. We observe that by expanding $K_{t}$ and $R$ in Equation˜4, the update equation can be reformulated as:
@@ 4.1 Embedding Misalignment in Effective Massive Editing
where the update $\Delta$ is determined by the aggregated residual and key embeddings across all edited facts. As the number of edits increases, solving Equation˜6 is more likely to cause reconstruction loss for individual knowledge items due to the embedding misalignment between the residual and key embeddings. This eventually leads to degraded editing performance.
@@
To formalize the concept of embedding misalignment, we define two key requirements for the desired update $\Delta$: (1) The update should preserve the existing knowledge, expressed as $\Delta C_{p}=0$. (2) The update should ensure lossless reconstruction for each individual fact, formulated as $\Delta k_{i}=r_{i}$, where $\Delta$ is computed while considering all target facts. Incorporating (1), an ideal $\Delta$ that meets (2) implies:
$\Delta\left(C_{p}+\sum_{i=1}^{N_{t}}k_{i}k_{i}^{T}\right)=\sum_{i=1}^{N_{t}}r_{i}k_{i}^{T}\quad\rightarrow\quad\Delta k_{i}=r_{i}\quad\text{for }i=1,2,\ldots,N_{t}$ (7)
@@
However, the validity of Equation˜7 is intuitively affected by the degree of misalignment between the residual and key embedding of different facts. We then define embedding misalignment:
Definition 1 (Embedding Misalignment). Given $N$ knowledge items, let each item $i$ be associated with a residual embedding $r_{i}$ and a key embedding $k_{i}$. We define the embedding misalignment of item $i$ as the structural similarity between the pairwise relations of its residual embedding and those of its key embedding. Formally, consider the distributions
$P_{r}^{(i)}=\{\,cos(r_{i},r_{j})\mid j\neq i\,\},\qquad P_{k}^{(i)}=\{\,cos(k_{i},k_{j})\mid j\neq i\,\},$ (8)
where $cos(\cdot,\cdot)$ is the cosine similarity. The $i$th misalignment score is quantified by the KL divergence:
$\mathcal{A}(i)=\mathrm{KL}\!\left(P_{r}^{(i)}\;\|\;P_{k}^{(i)}\right).$ (9)
@@
We now formalize the connection between embedding misalignment and the editing performance of a specific knowledge item $i$ under massive editing. Specifically, we quantify the degree to which Equation˜7 is established by analyzing the reconstruction loss $e_{i}=\Delta k_{i}-r_{i}$ for each knowledge item. This relationship is formalized in the following theorem:
Theorem 1. Let $\Delta$ be the closed-form solution satisfying $\Delta\sum_{i}k_{i}k_{i}^{\top}=\sum_{i}r_{i}k_{i}^{\top}$, and define the reconstruction residual of item $i$ as $e_{i}=\Delta k_{i}-r_{i}$. Then we can expand
$e_{i}=\sum_{j=1}^{N}\beta_{ij}r_{j}-r_{i},\qquad\beta_{ij}:=k_{j}^{\top}\!\Big(\sum_{\ell=1}^{N}k_{\ell}k_{\ell}^{\top}\Big)^{\!-1}\!k_{i}$ (10)
and its norm is bounded by the misalignment between the neighborhood structures of $r_{i}$ and $k_{i}$:
$\|e_{i}\|\;\leq\;C_{i}\sqrt{\tfrac{1}{2}\,\mathcal{A}(i)}\;+\;|\beta_{ii}|\,\|r_{i}\|\;+\;\|\varepsilon_{i}\|,$ (11)
@@
This result demonstrates how embedding misalignment impacts the editing performance of individual knowledge items under massive editing. Specifically, stronger misalignment among knowledge items leads to increased individual reconstruction loss, ultimately reducing the overall effectiveness of massive editing. The complete proof is provided in Appendix˜B.
Table 1: Editing performance with varying numbers of edited facts on LLaMA2-7B and Deepseek-7B.
Model LLaMA2-7B Deepseek-7B
No. of Edited Facts 200 500 1000 200 500 1000
Editing Efficiency(%) 98.5 90.0 86.8 99.5 98.6 97.8
$\sum_{i}\mathcal{A}(i)$ 79 243 554 68 223 562
@@
Empirical Study. Motivated by the above analysis, we hypothesize that the failure of massive editing stems from misalignment between the embeddings of different knowledge items. To test this hypothesis, we edit 200, 500, and 1,000 facts from the CounterFact dataset (Meng et al., 2022) using MEMIT (Meng et al., 2023) on LLaMA2-7B (Touvron et al., 2023a) and Deepseek-7B (Bi et al., 2024). We then evaluate the editing accuracy of these items when no prefix is added to the edited query. Embedding misalignment is quantified using the misalignment score defined in Equation˜9.
As shown in Table˜1, the overall editing accuracy of both models decreases as more facts are edited, accompanied by a clear increase in embedding misalignment. For example, on LLaMA2-7B, the accuracy drops from 98.5% to 86.8% as the number of edited facts grows from 200 to 1,000, while the misalignment score rises from 79 to 554. These results provide further evidence for our theorem that embedding misalignment leads to degraded editing performance.
4.2 Impact of Embedding Misalignment on Editing Robustness
We investigate how embedding misalignment affects robustness in massive editing along two dimensions: (i) long-prefix perturbations and (ii) simultaneous edits of samples sharing the same subject. Based on our theoretical and empirical analysis, we derive two corollaries to characterize these effects and validate them with controlled experiments.
Corollary 1. Long prefixes exacerbate embedding misalignment issues under massive editing, leading to degraded editing performance when edited facts are evaluated with descriptive prefixes.